# Data-driven Virtual Test-bed of the Blown Powder Directed Energy Deposition Process

### v1

@@ Abstract
Digital twins in manufacturing serve as a crucial bridge between the industrial age and the digital age, offering immense value. Current additive manufacturing processes are able to generate vast amounts of in-process data, which, when effectively ingested, can be transformed into insightful decisions. Data-driven methods from reduced order modeling and system identification are particularly promising in managing this data deluge. This study focuses on Laser Powder Directed Energy Deposition (LP-DED) equipped with in-situ process measurements to develop a compact virtual test-bed. This test-bed can accurately ingest arbitrary process inputs and report in-process observables as outputs. This virtual test-bed is derived using Dynamic Mode Decomposition with Control (DMDc) and is coupled with uncertainty quantification techniques to ensure robust predictions.
@@ 1 Introduction
Additive manufacturing, a relatively young set of practices, has garnered significant attention and promise over the past few decades. This promise lies in its ability to enable designers and engineers to create better, lighter, and stronger products. However, unlike traditional manufacturing methods, additive manufacturing introduces new challenges, particularly significant uncertainties between input process parameters and the final product [33, 39]. These uncertainties can range from variations in material performance to shape defects and geometrical non-conformance [24, 38]. While these issues affect different additive manufacturing techniques to varying degrees, this study focuses on directed energy deposition.
Directed energy deposition (DED), also known as laser metal deposition (LMD), direct metal deposition (DMD), and colloquially as LENS®, is one of the seven archetypal forms of additive manufacturing [13, 7, 1]. In DED, metal powder is delivered to a substrate where a heat source, typically a laser, simultaneously induces melting, thereby depositing material. The preferred machine arrangement in DED is still under debate, but for this study, we use a laser heat source with the powder feedstock delivered coaxially to the laser beam through an annulus nozzle.
To better understand the link between input process parameters and the final product, on-machine monitoring (OMM) is employed during the process. OMM involves using sensors to record the in-process state. Common sensors measure the size, temperature, and light emission intensity of the molten pool, but these represent only a small subset of available sensor modalities [15, 11]. In this study, the OMM suite on the DED machine includes coaxial measurement of melt pool size and temperature, as well as an off-axis camera that measures the working distance between the nozzle and melt pool centroid. Additional measurements are summarized in Table 1. While having this in-process information is valuable, our focus is on managing and generating impactful insights from the data that can be applied to previously unseen input/actuation.
On-Machine Monitoring (OMM) Capability
Process Monitoring Sensors (states) Process Inputs Environment Monitoring Sensors
@@
Laser Power Feedback (Watts) Powder Hopper Disc Speed (RPM)
@@
System models are indispensable tools in engineering, science, and various other fields, providing a structured approach to understanding, predicting, and optimizing the behavior of complex systems. These models form the foundation of what are known as digital twins [8]. Digital twins are virtual replicas of physical systems that are used to simulate, analyze, and optimize real-world processes. We have adopted the term "virtual test-bed" to specifically highlight not only the replication of processes but also the extension to unseen inputs and actuation. Hartmann et al. have summarized the current state of digital twins/system models in the context of additive manufacturing (AM), noting that achieving an accurate system model with relatively low computational cost remains a significant milestone yet to be reached [18].
These models can be broadly categorized into physical (white-box), hybrid (gray-box), and data-driven (black-box), each with its unique advantages and applications. In DED, both Mukherjee and Khairallah et al. stressed the need for accurate, physical simulation models [30, 25]. Other physical models have also been presented by Ertay et al. and Weisz-Patrault [12, 42]. To contrast, Gaikwad et al. demonstrated a hybrid, physical and data-driven model that combined physical simulation with in-situ sensor data and showed that the predictive capacity increased through this combination [15].
Our work presented here is a purely data-driven model which relies entirely on empirical data to make predictions without any explicit knowledge of the system’s internal workings. These models use algorithms to identify patterns and relationships within the data, making them highly versatile and capable of handling complex, nonlinear systems. They do, however, require large amounts of high-quality data for training and can sometimes lack interpretability, making it challenging to understand how predictions are made. Within this data-driven paradigm, two key sub-branches are system identification and reduced order modeling (ROM). Among these techniques, dynamic mode decomposition (DMD) introduced by Schmid and its variants, such as dynamic mode decomposition with control (DMDc), have gained prominence for their ability to extract low-order models from high-dimensional data [34, 31].
ROMs aim to simplify complex systems by reducing the number of degrees of freedom while retaining essential dynamic characteristics. Techniques such as proper orthogonal decomposition (POD) and DMD are commonly used [3, 34]. POD identifies the most energetic modes in a system, while DMD decomposes the system into modes that capture both spatial and temporal dynamics. Autoregressive models, which predict future states based on past observations, are also widely used. System identification involves developing mathematical models of dynamic systems based on measured data [28]. Techniques in this area include output-error methods, which minimize the difference between observed and predicted outputs, and Hammerstein-Wiener models, which combine linear dynamic models with static nonlinearities. Subspace identification methods use state-space representations to identify system dynamics from input-output data. Other traditional methods, such as the Eigensystem Realization Algorithm (ERA) and Observer Kalman Filter Identification (OKID), use connections between measurements and observability and controllability matrices from control theory [22, 23]. However, these methods focus on single impulse-response measurements, which, in an additive manufacturing context, would require averaging across many derived system matrices to achieve the same consensus as DMD/DMDc.
In recent years, machine learning techniques have become integral to system identification and ROM. Neural networks, which can approximate complex nonlinear functions, are increasingly used to model system dynamics. Long Short-Term Memory (LSTM) networks, a type of recurrent neural network, are particularly effective for time-series prediction due to their ability to capture long-term dependencies. AutoEncoders, which learn efficient representations of data, are used for dimensionality reduction and feature extraction [16]. Notable recent advancements in the field include Sparse Identification of Nonlinear Dynamics (SINDy) and Latent Space Dynamics Identification (LaSDI) [4, 14]. SINDy identifies governing equations of a system by promoting sparsity in the model, making it interpretable and efficient. LaSDI leverages latent space representations to identify dynamics, enabling the modeling of complex systems with reduced computational cost.
In this work, we employ DMDc, a derivative of DMD. DMDc extends the DMD framework by incorporating control inputs, allowing it to assess the influence of these inputs on the system’s state [31]. This is particularly useful in systems where control actions play a significant role in the dynamics. Traditionally, handling multiple control input points required constructing multiple DMD models and interpolating between state-transition matrices [40]. This approach can be cumbersome and computationally expensive. DMDc simplifies this process by factoring out the confounding effects of control inputs. It constructs a single model that can handle multiple control inputs, making it more efficient and easier to implement. This is achieved by augmenting the state-space representation with control inputs, allowing the model to capture the influence of these inputs on the system dynamics.
A detailed description of the DMDc model formulation is provided in Section 2.1, but it is important to highlight some key features beforehand. DMDc is closely related to linear state-space models and shares the primary assumption that the system is Markovian in nature. This means that the future state of the system depends only on the current state and control inputs, not on past states. Consequently, careful consideration of both spatial and temporal discretization is required in the analysis to ensure accurate modeling.
The literature on dynamic mode decomposition and its variants, such as DMDc, reveals significant advancements in data-driven modeling of complex systems [35]. DMDc offers a robust framework for incorporating control inputs and disambiguate system dynamics from actuation effects, while modal decomposition techniques like POD and Balanced POD [43] provide valuable insights into fluid flows and other high-dimensional systems. However, challenges related to noise, nonlinearity, and computational expense remain [20, 19, 44]. Future research should focus on refining these methods to enhance their robustness and applicability across various fields, ultimately improving our ability to model and understand complex systems. This study leverages DMDc to model the complex dynamics of directed energy deposition (DED) in additive manufacturing. By incorporating control inputs into the DMD framework, we aim to better understand the link between input process parameters and the final product, ultimately improving the reliability and performance of DED processes, which are significantly actuation-dependent.
2 Methods and Materials
The dataset utilized in this study comprises 26 individual experiments, each designed to investigate various process parameters of the DED process using Inconel 625. The experiments were conducted under controlled conditions to ensure the reliability and reproducibility of the data. For the purpose of model training and validation, the data is split such that 25 experiments are used for training the model, and one experiment is reserved for testing. This approach ensures that the model is evaluated on unseen data, providing a robust assessment of its predictive capabilities. Data is generated during the DED process at a user-defined rate, which can range from 10 to 500 Hz, with a default rate of 100 Hz. This high-frequency data collection captures the dynamic nature of the process, providing detailed insights into the behavior of the system.
@@ 2 Methods and Materials
The proposed virtual test-bed workflow begins with the creation of a G-code program. G-code is the standard programming language used to control CNC machines, including those used in DED processes. It dictates the machine’s movements, laser modulation, and other critical parameters necessary for part fabrication. The G-code program is converted into a time series representation of all the inputs by the machine controller. This conversion translates the discrete commands into a continuous time series that can be analyzed and processed. The time series data undergoes a transformation where the inputs are standardized to have a zero mean and unit standard deviation. This step is crucial for ensuring that the data is on a common scale, which facilitates more effective modeling. The transformation parameters are saved for later use in inverse transformation. Although not strictly necessary, the authors have found that standardizing and normalizing the process inputs yield better results. Standardization is advantageous for mathematical reasons, primarily because it unitizes the input space, making the data more amenable to analysis and modeling.
The standardized inputs are fed into a state-space model, which iteratively predicts the observables or sensor measurements for the next time step until all inputs are processed. The state-space model is derived using DMDc, which is performed during a separate training stage of the pipeline. The predictions are then inverse transformed back into their original space, ensuring that the outputs are in the same units as the original inputs. These predictions are combined with uncertainty information to provide a comprehensive understanding of the model’s confidence in its predictions. The uncertainty information from previous builds is gathered separately with OMM data from various builds forming a cache of uncertainty information to be applied later. This cache provides valuable insights into the variability and reliability of the process, enhancing the robustness of the model’s predictions. The authors believe that this cache improves with the ingestion of more data, similar to the DMDc state-space model, thereby continuously refining the model’s accuracy and reliability. The final output of the workflow includes both a time series and a geometric digital representation of the final object. This dual output provides a detailed view of the process dynamics and the resulting part geometry.
In summary, the methods and materials described in this study provide a comprehensive framework for modeling and predicting the behavior of the DED process using Inconel 625. The integration of advanced data transformation techniques, state-space modeling, and uncertainty quantification ensures that the proposed workflow is both robust and reliable, paving the way for more effective control and optimization of additive manufacturing processes.
@@
2.1 Base Mathematical Formulation
At a fundamental level, DMD analyzes the relationship between pairs of measurements from a dynamical system. The measurements, $x_{t}$ and $x_{t+1}$, where $t$ denotes the temporal iteration from a discrete dynamical system, are assumed to be approximately related by a linear operator:
$x_{t+1}\approx\textbf{A}x_{t},$ (1)
where $x\in\mathbb{R}^{n}$ and $\mathbf{A}\in\mathbb{R}^{n\times n}$. This approximation is presumed to hold for all pairs of measurements. The subsequent discussion of DMD will focus on determining a best-fit solution for the operator $\mathbf{A}$ across all measurement pairs. It is important to note that the relationship in Equation 1 does not need to hold exactly, as is typical with most numerical methods.
DMDc extends Equation 1 to encompass a more comprehensive state-space representation. The state-space representation of a linear dynamical system is given by:
$x_{t+1}=\textbf{A}x_{t}+\textbf{B}u_{t}$ (2)
$y_{t}=\textbf{C}x_{t}+\textbf{D}u_{t}$
Assuming no direct feed-through from inputs to measurements ($\mathbf{D}=0$) and setting $\mathbf{C}$ to the identity matrix ($\mathbf{C}=\mathbf{I}$), we imply that there is no access to internal states ($y_{t}=x_{t}$). This assumption aligns with the data-driven hypothesis and the gappy measurement problem [17]. However, it is noteworthy that the development and previous applications of DMD utilized fluid dynamics numerical simulations where full-state access was available. The more general case described here employs measurements of a system, whether historical, numerical, or experimental, to construct the data matrix. A control-theoretic perspective would frame Equation 1 with a measurement function. A linear measurement function $x=\mathbf{C}x$ suggests an underlying system involving the variable $x$. In Section 8, we demonstrate how this measurement function can take different forms. Substituting into Equation 2 yields:
$y_{t+1}=\textbf{A}y_{t}+\textbf{B}u_{t}$ (3)
which relates the system dynamics solely as a function of input-output pairs and fits into the DMDc framework. From Equation 2, given the process inputs and recorded measurements, $\mathbf{A}$ and $\mathbf{B}$ are unknown, with $y_{t}$ and $u_{t}$ being replaced by snapshot matrices as shown in Equation 4:
To draw equivalence with standard DMD, Proctor et al. introduces two new variables, $\mathbf{G}$ and $\Omega$, to combine column-wise the unknowns and the measurements and process inputs (Equation 5) [31]:
$\textbf{G}=[A|B]=\textbf{Y}_{t+1}\Omega^{\dagger}\approx\textbf{Y}_{t+1}\times\textrm{SVD}\biggl{(}\Bigl{[}\frac{\textbf{Y}_{t}}{\textbf{U}_{t}}\Bigr{]}\biggr{)}^{\dagger}$ (6)
The $\dagger$ symbol represents the Moore-Penrose pseudoinverse. Canonically, the singular value decomposition (SVD) uses $U$ and $V$ to represent the unitary matrices. However, to avoid confusion with the process input snapshot matrix $\mathbf{U}$, the SVD will be taken as:
Applying the SVD as previously defined to Equation 6, $\mathbf{G}$ can be solved for both $\mathbf{A}$ and $\mathbf{B}$ as:
$\textbf{A}=\textbf{Y}_{t+1}\zeta\Sigma^{-1}\eta_{Y_{t}}^{*}$ (8)
$\textbf{B}=\textbf{Y}_{t+1}\zeta\Sigma^{-1}\eta_{U_{t}}^{*}$
where $\eta_{Y_{t}}^{*}$ and $\eta_{U_{t}}^{*}$ are the measurement and process input components of the decomposition, respectively. Note that a reduced dimensionality SVD as well as projection matrices typical of DMD/DMDc are not used in this formulation. The measurement space is extremely limited (consistent with the gappy type problem), and therefore all measurements are considered linearly independent in the context of a state-space representation.
2.2 Control Theoretic Lemmas
There are two equally important aspects of the collected experimental data that contribute to the success of the DMDc representation of the true system: the density of time snapshots and the induced dynamics observed during the experiment.
@@ 2.2 Control Theoretic Lemmas
This study examines the problem from two perspectives: inter-signal and intra-signal analysis. Inter-signal analysis, commonly referred to as feature selection in the fields of statistics, machine learning, and data science, is a critical aspect of this study [5]. The three control theoretic lemmas discussed earlier primarily fall under the umbrella of feature selection. Feature selection encompasses a set of tools and practices designed to address degeneracies in the data, thereby mitigating their adverse effects on model performance. Two prominent forms of data degeneracy are the curse of dimensionality and multicollinearity, both of which manifest similarly in regression analysis. In the limit, both the curse of dimensionality and multicollinearity tend to produce regression coefficients with little to no inter-feature sensitivity. Common techniques for addressing these issues include correlation analysis, matrix rank assessment, chi-squared tests, mutual information, recursive feature elimination, and, as utilized in this study, the variance inflation factor (VIF) method [36].
Due to the limited number of observables available, they were not subjected to any feature selection process. However, the process inputs were analyzed for multicollinearity using the VIF method. The VIF method allows for the exploration and hierarchy of inputs that produce a multicollinearity condition, whereas matrix rank assessment merely indicates the existence of such a condition. The algorithm for performing the VIF method involves selecting an individual feature and then forming a regression equation using the remaining features. The coefficients of this regression are then solved (see Equation 9). The VIF for that feature is calculated as shown in Equation 10, where ($R^{2}$) is the coefficient of determination from the solved regression. This process is repeated for each individual feature. Empirically, VIF values greater than 10 indicate that the given feature exhibits a high degree of multicollinearity, while VIF values less than 5 suggest that the feature is independent of all other features [36]. Our feature selection algorithm involves removing features with VIF values greater than 10 and iterating through the VIF method until all features have VIF values less than 5. Additionally, we assess the traditional matrix rank at every iteration to ensure that the process is not erroneously halted.
@@
Intra-signal/feature analysis, often referenced in the context of techniques such as SMOTE (Synthetic Minority Over-sampling Technique) and Intelligent Feed Forward (IFF), is best understood through the lens of a classification problem [6, 10]. Consider a scenario where a single feature is categorized into six distinct classes. If there is an imbalance in the number of examples for each class, the resulting model is likely to exhibit bias. This imbalance can lead to the under-representation or even elimination of classes with fewer observations, driven by mechanisms related to the loss function or inherent uncertainty in the model. Extending this analogy to a continuous case, a balanced feature would be one that approximates a uniform distribution over its given range.
@@
Our formulation and attempt at characterizing the "intra-feature rank" involves measuring the distance between distributions, which can be effectively accomplished using the Wasserstein distance, also known as the Earth Mover’s Distance. The Wasserstein distance quantifies the amount of work required to transform one distribution into another, providing a robust metric for assessing the similarity between distributions. By employing this metric, we can rigorously evaluate the balance of a feature, ensuring that it does not introduce bias into the model due to distributional imbalances.
@@
In summary, this study employs a comprehensive approach to data analysis, addressing both inter-signal and intra-signal aspects. The inter-signal analysis, through feature selection and multicollinearity assessment using the VIF method, ensures that the data set is robust and free from degeneracies that could impair model performance. The intra-signal analysis ensures that the temporal characteristics of the data are accurately captured, providing a solid foundation for constructing reliable and accurate models of the underlying physical processes.

### v2

# Harnessing On-Machine Metrology Data for Prints with a Surrogate Model for Laser Powder Directed Energy Deposition

@@ Abstract
In this study, we leverage the massive amount of multi-modal on-machine metrology data generated from Laser Powder Directed Energy Deposition (LP-DED) to construct a comprehensive surrogate model of the 3D printing process. By employing Dynamic Mode Decomposition with Control (DMDc), a data-driven technique, we capture the complex physics inherent in this extensive dataset. This physics-based surrogate model emphasizes thermodynamically significant quantities, enabling us to accurately predict key process outcomes. The model ingests 21 process parameters, including laser power, scan rate, and position, while providing outputs such as melt pool temperature, melt pool size, and other essential observables. Furthermore, it incorporates uncertainty quantification to provide bounds on these predictions, enhancing reliability and confidence in the results. We then deploy the surrogate model on a new, unseen part and monitor the printing process as validation of the method. Our experimental results demonstrate that the predictions align with actual measurements with high accuracy, confirming the effectiveness of our approach. This methodology not only facilitates real-time predictions but also operates at process-relevant speeds, establishing a basis for implementing feedback control in LP-DED.
@@ 1 Introduction
Additive manufacturing has gained significant attention for its potential to create innovative, lightweight, and strong products. However, it presents unique challenges, particularly uncertainties in how input parameters affect the final product, including material performance variations and geometrical defects [26, 32, 18, 31]. This study focuses on laser powder directed energy deposition (LP-DED), an additive manufacturing technique where, in this specific case, metal powder is coaxially delivered and melted using a laser [11, 6, 1].
Modeling the LP-DED process involves a comprehensive understanding of the thermal, fluid, and mechanical interactions that occur during the additive manufacturing process. These models typically encompass heat transfer phenomena, fluid dynamics and solidification behavior of the molten pool. In DED, Mukherjee stressed the need for accurate, physical simulation models [22]. Other physical models have also been presented by Ertay et al. and Weisz-Patrault [10, 36]. The development of these physical models is crucial for the creation of digital twins, an umbrella term, which serve as virtual replicas of the LP-DED process [7]. Digital twins integrate real-time data from the manufacturing environment with physical models, enabling a dynamic simulation that reflects the actual operating conditions [13, 23]. By combining physical models with OMM sensor data, as demonstrated by Gaikwad et al., the predictive capacity of the digital twin can be significantly enhanced [12].
Hartmann et al. have summarized the current state of digital twins in the context of additive manufacturing (AM), noting that achieving an accurate system model at a relatively low computational cost remains a significant milestone yet to be reached [15]. We concur with this assessment, noting that accurate physical models are challenged by the multi-scale nature of AM, while machine learning (ML) or data-driven models often struggle with generalization. Furthermore, reviews by Wang et al. and Chen et al. on data-driven model and in-situ monitoring, respectively, indicate that a very small proportion of models in the literature are "computationally efficient" [34, 5]. The models in the literature can be broadly grouped into four classes: Linear, Gaussian Process (GP), Tree, and Multi-Layer Perceptron (MLP). Wang reports on 70 models for laser AM, both LPBF and DED, where MLPs enjoy the majority at 51.4% of those reported. Similarly, Chen reports that MLP models occupy 63.2% in DED, and for good reason as MLPs, especially in other fields, have seen fantastic success when applied on those problems. MLPs do scale well with data size, however, they tend to shift the computational cost away from prediction and into training the model. The next largest class, GP-based models, accounts for 34.7% as per Wang. GPs have advantageous properties, including the provision of uncertainty bounds and ease of sampling [25]. GP models have been utilized in numerous AM related regression problems with varying levels of success, but their primary drawback is the inherent computational complexity, which is ($O(n^{3})$), imposing a practical limit on the number of data points that can be fit [2, 21]. Linear and tree-based models represent the smallest fraction in the AM literature. Linear models, often perceived as overly simplistic for capturing the complex behavior in laser AM, are computationally efficient for both training and prediction. We are investigating whether innovative linear models can accurately represent the DED process with low computational cost.
Our study introduces a surrogate state-space model of the LP-DED process, where the state transition and exogenous input operators are derived using dynamic mode decomposition with control (DMDc). Falling under the aforementioned class of linear models, DMDc is a variant of the original dynamic mode decomposition algorithm introduced by Schmid, designed for extracting low-order models from high-dimensional data in the presence of exogenous inputs [27, 24, 28]. This study aims to accurately connect process inputs to OMM observables, incorporating uncertainty bounds. Recently, Hermann et al. proposed a similar workflow to ours, where their model processes input parameters to predict deposition geometry, ultimately to optimize tool-path planning and the manufacturing of complete DED parts [16]. Their primary assertion is that entire builds can be predicted as the cumulative result of multiple single-tracks, which, in our experience, is valid but not entirely comprehensive. We contend that enhancing the reliability and repeatability of the DED process at scale necessitates a more detailed perspective, one that considers both spatial and temporal components of process inputs and their effects on observables.
We demonstrate that with careful consideration of the training data and the incorporation of uncertainty bounds, a fast linear state-space model can serve as an accurate surrogate for the LP-DED process. Our goal is to enhance the reliability and performance of DED processes by better linking input parameters to final outcomes, thereby addressing a critical gap in current additive manufacturing research.
2 Methods and Materials
For this study, the OMM suite on the DED machine includes coaxial measurement of melt pool size and temperature, as well as an off-axis camera that measures the working distance between the nozzle and melt pool with additional measurements are summarized in Table 1. Additional details on the machine and OMM suite available in [17]. The dataset utilized comprises 26 individual experiments, each designed to investigate various process parameters of the DED process using Inconel 625. The experiments were conducted under controlled conditions to ensure the reliability and reproducibility of the data. For the purpose of model training and validation, a "Leave p Out Cross Validation" (LpOCV) strategy was used to split the OMM dataset into train and testing datasets with $p=3$. This approach ensures that the model is evaluated on unseen data, providing a honest assessment of its predictive capabilities. Data is generated during the DED process at a user-defined rate, which can range from 10 to 500 Hz, with a default rate of 100 Hz. This high-frequency data collection captures the dynamic nature of the process, providing detailed insights into the behavior of the system.
On-Machine Monitoring (OMM)
Capability
Process Monitoring Sensors Process Inputs Environment Monitoring
(states) Sensors
@@ 2 Methods and Materials
Laser Power Feedback (Watts) A Position (deg)
C Position (deg)
Powder Hopper Disc Speed (RPM))
@@ C Position (deg)
Laser Power On Time (s)
NC Line Number (int)
Part Distance Traveled (mm)
@@
The proposed workflow begins with a G-code program. G-code is the standard programming language used to control CNC machines, including those used in DED processes and it dictates the machine’s movements, laser modulation, and other critical parameters necessary for part fabrication. The G-code program is converted into a time series representation of all the inputs by the machine controller that translates the discrete commands into a continuous time series that can be analyzed and processed. This time series data undergoes a transformation where the inputs are standardized to have a zero mean and unit standard deviation. Although not strictly necessary, the authors have found that standardizing and normalizing the process inputs yielded better results. The transformation parameters are saved for later use in inverse transformation.
The standardized inputs are fed into a state-space model, which iteratively predicts the observables or sensor measurements for the next time step until all inputs are processed. The state-space model is derived using DMDc, which is performed during a separate training stage of the pipeline. A detailed description of the DMDc model formulation is provided in section 2.3, but it is important to highlight some key features beforehand. DMDc is closely related to linear state-space models and shares the primary assumption that the system is Markovian in nature. This means that the future state of the system depends only on the current state and control inputs, not on past states. Consequently, careful consideration of both spatial and temporal discretization is required in the analysis to ensure accurate modeling.
The predictions are subsequently inverse transformed to their original scale and variance, ensuring that the outputs are consistent with the units of the original inputs. These predictions are integrated with uncertainty information to provide insight into the model’s confidence in its predictions. The uncertainty information is specifically derived from the LpOCV DMDc training process, where error levels and variability statistics are assessed, allowing for the inclusion of minimum and maximum error envelopes in the state-space predictions. The final output of the workflow comprises both a time series and a geometric digital representation of the final object. This dual output offers a comprehensive view of the process dynamics and the resulting part geometry.
@@
2.1 The Dataset
The raw dataset, detailed in Section 2, comprises of 26 DED experiments involving simple square prisms, conducted as part of a process parameter development study for Inconel 625. This dataset includes 49 columns (features) and 1,706,677 rows (samples). Most features are listed in Table 1, with some omitted due to redundancy (e.g., Powder Hopper Disc Speed 1, 2, etc.). The dataset was examined for missing data (NaNs) and multicollinearity (repetitive columns). Two columns, PMD ($g/mm^{2}$) and Powder Feeder 2 Mass Flow Rate ($g/min$), initially contained NaN values and were excluded, as these features—recorded by the machine’s DAQ—require additional calibration steps that were not performed and their exclusion was therefore expected. Beyond these features, no other columns contained NaN values. Any corrections for NaNs would have been addressed on a case-by-case basis, given the time-series nature of the analysis. Following this, the dataset was examined for inactive features (e.g., inactive powder feeders), which were also removed. Multicollinearity was addressed using a feature selection algorithm based on the Variance Inflation Factor (VIF), reducing the dataset to 21 features. The VIF algorithm was applied specifically to process input features. Initially, the 49 features were stripped of any state-related or inactive features, reducing the dataset to 28 features before applying the VIF algorithm. The test:train split followed a "Leave 3 Out Cross-Validation" methodology, where three experiments were selected to represent the test dataset, and the remaining experiments were used for training. Notably, the dataset was not standardized prior to the split to avoid introducing potential data bias.
2.2 OMM Feature Selection and Curation
There are two equally important aspects of the collected experimental data that contribute to the success of a DMDc representation of the true system: the density of time snapshots and the induced dynamics observed during the experiment.
@@ 2.2 OMM Feature Selection and Curation
Our interpretation of the above control theoretic lemmas (List 2.2) is that no degeneracies exist in either the states or inputs separately or in combination. For this study we assume no degeneracy exists in the states, we test for degeneracies in the inputs, and as a consequence we must also assume there is no state feedback in our data.
This study examines the problem from two perspectives: inter-signal and intra-signal analysis. Inter-signal analysis, commonly referred to as feature selection in the fields of statistics, machine learning, and data science, is a critical aspect of this study [3]. The three control theoretic lemmas discussed earlier primarily fall under the umbrella of feature selection. Feature selection encompasses a set of tools and practices designed to address degeneracies in the data, thereby mitigating their adverse effects on model performance. Two prominent forms of data degeneracy are the curse of dimensionality and multicollinearity, both of which manifest similarly in regression analysis. In the limit, both the curse of dimensionality and multicollinearity tend to produce regression coefficients with little to no inter-feature sensitivity. Common techniques for addressing these issues include correlation analysis, matrix rank assessment, chi-squared tests, mutual information, and recursive feature elimination among others. Matrix rank assessment, crucial to the control theoretic lemmas, we find as a necessary but not sufficient condition for characterizing inter-signal degeneracy and therefore have coupled it to another feature selection method, the variance inflation factor (VIF) method [29].
Due to the limited number of observables available, they were not subjected to any feature selection process, hence the full rank assumption earlier. However, the process inputs were analyzed for multicollinearity using a matrix rank assessment along with the VIF method. The VIF method allows for the exploration and hierarchy of inputs that produce a multicollinearity condition, whereas matrix rank assessment merely indicates the existence of such a condition. The algorithm for performing the VIF method involves selecting an individual feature and then forming a regression equation using the remaining features. The coefficients of this regression are then solved (see (1)). The VIF for that feature is calculated as shown in (2), where ($R^{2}$) is the coefficient of determination from the solved regression. This process is repeated for each individual feature. Empirically, VIF values greater than 10 indicate that the given feature exhibits a high degree of multicollinearity, while VIF values less than 5 suggest that the feature is independent of all other features [29]. Our feature selection algorithm involves removing features with VIF values greater than 10 and iterating through the VIF method until all features have VIF values less than 5. The matrix rank assessment performed at every iteration ensures that the process is not erroneously halted.
@@
Intra-signal/feature analysis, often referenced in the context of techniques such as SMOTE (Synthetic Minority Over-sampling Technique) and Intelligent Feed Forward (IFF), is best understood through the lens of a classification problem [4, 9]. As an illustrative example, consider a scenario where a single feature is categorized into six distinct classes. If there is an imbalance in the number of examples for each class, the resulting model is likely to exhibit bias. This imbalance can lead to the under-representation or even elimination of classes with fewer observations, driven by mechanisms related to the loss function or inherent uncertainty in the model. Extending this analogy to a continuous case, a balanced feature would be one that approximates a uniform distribution over its given range and we therefore utilize it as a theoretical benchmark for a given signal.
@@
Our formulation and characterization of the "intra-feature rank" involves measuring the distance between distributions, which can be effectively accomplished using the Wasserstein distance, also known as the Earth Mover’s Distance. The Wasserstein distance quantifies the amount of work required to transform one distribution into another, providing a flexible metric for assessing the similarity between distributions. It is recognized as a true metric, and by employing it, we can rigorously evaluate the similarity of the data being ingested by the model, as well as assess the dynamic quality of the features.
@@
In summary, this study employs a comprehensive approach to data analysis, addressing both inter-signal and intra-signal aspects. The inter-signal analysis, through feature selection and multicollinearity assessment using a matrix rank assessment and the VIF method, ensures that the data set is robust and free from degeneracies that could impair model performance. The intra-signal analysis ensures that individual features are distributionally similar both of which provide a solid foundation for constructing reliable and accurate models of the underlying physical processes.
2.3 State-Space Model: DMDc Formulation
At a fundamental level, DMD analyzes the relationship between pairs of measurements from a dynamical system. Likely the most basic version of a dynamical system is summarized in (3) below where $x_{t}$ is the current state of the system, $x_{t+1}$ is a future state with these states approximately related by a linear operator, $\mathbf{A}$. DMD looks to solve for the linear operator, $\mathbf{A}$, which is sometimes called the state transition matrix.
$x_{t+1}\approx\textbf{A}x_{t},$ (3)
It is important to note that the relationship in (3) does not need to hold exactly a fact that DMD leverages in that equality holds in the "least-squares" sense. We will utilize a more different form of (3) to include the effects of actuation and measurement. Our discrete, state-space representation of the linear dynamical system of interest will be given by:
$x_{t+1}=\textbf{A}x_{t}+\textbf{B}u_{t}$ (4)
$y_{t}=\textbf{C}x_{t}+\textbf{D}u_{t}$
Where again we have the state vector $x\in\mathbb{R}^{n}$ and the linear operator $\mathbf{A}\in\mathbb{R}^{n\times n}$. Additionally, the input (or control) vector, $u\in\mathbb{R}^{p}$, and measurement vector, $y\in\mathbb{R}^{q}$, along with the input matrix, $\mathbf{B}\in\mathbb{R}^{n\times p}$, state-to-measurement matrix, $\mathbf{C}\in\mathbb{R}^{q\times n}$, and feed-through matrix, $\mathbf{D}\in\mathbb{R}^{q\times p}$.
It is of note that the development and many previous applications of DMD utilized fluid dynamics numerical simulations where full-state access was available [27, 33, 28]. In this data-driven paradigm with only a limited amount of sensors capable of recording the system, access to the full-state is not possible. Therefore, changes to the state-space equations are required for the problem at hand. Interestingly, limited state measurements are an issue faced in multiple domains and has been termed the gappy measurement problem [14]. There are no current strategies which address the gappy measurement holistically and must addressed in an individualized manner. To that end, we assume no direct feed-through from inputs to measurements ($\mathbf{D}=0$), and by setting $\mathbf{C}$ to the identity matrix ($\mathbf{C}=\mathbf{I}$) we imply that there is no access to internal states ($y_{t}=x_{t}$). Substituting these conditions into (4) yields:
$y_{t+1}=\textbf{A}y_{t}+\textbf{B}u_{t}$ (5)
which relates the system dynamics solely as a function of input-measurement pairs. Examining the equation in (5), we are given the process inputs and recorded measurements thought the collected OMM data and therefore $\mathbf{A}$ and $\mathbf{B}$ are the unknowns that must be solved. To relate (5) with (3), Proctor et al. suggests that the snapshot matrices be modified slightly as shown in (6) and introduces two new variables, $\mathbf{G}$ and $\Omega$, to combine column-wise the unknowns and the measurements and process inputs (7) [24].
Now, solving for $\mathbf{G}$ we take the product of the measurement snapshot matrix, $\textbf{Y}_{t+1}$ and $\Omega$ inverse. Numerically the inverse is handled by the singular value decomposition (SVD) and in this case the SVD has additional mathematical advantages which allow us to solve for $\mathbf{A}$ and $\mathbf{B}$.
$\textbf{G}=[A|B]=\textbf{Y}_{t+1}\Omega^{\dagger}\approx\textbf{Y}_{t+1}\times\textrm{SVD}\biggl{(}\Bigl{[}\frac{\textbf{Y}_{t}}{\textbf{U}_{t}}\Bigr{]}\biggr{)}$ (8)
Canonically, the SVD uses $U$ and $V$ to represent the unitary matrices, however, to avoid confusion with the process input snapshot matrix $\mathbf{U}$, the SVD will be taken as:
Applying the SVD to (7) as shown in (8), $\mathbf{G}$ can be isolated and both $\mathbf{A}$ and $\mathbf{B}$ can be solved as:
$\textbf{A}=\textbf{Y}_{t+1}\zeta\Sigma^{-1}\eta_{Y_{t}}^{*}$ (10)
$\textbf{B}=\textbf{Y}_{t+1}\zeta\Sigma^{-1}\eta_{U_{t}}^{*}$
where $\eta_{Y_{t}}^{*}$ and $\eta_{U_{t}}^{*}$ are the measurement and process input components of the decomposition, respectively. Note that a reduced dimensionality SVD as well as projection matrices typical of DMD/DMDc are not used in this formulation due to the extremely limited measurement space (again consistent with a gappy-type measurement problem).
Lastly, for all DMDc computations of the system matrices, $\mathbf{A}$ and $\mathbf{B}$, using the OMM dataset we employ a "Leave p Out Cross Validation" (LpOCV) strategy with $p=3$. Computations then generally follow a format where 3 individual experiments are selected at random from the 26 overall experiments. Those 3 experiments serve as the test data with the remaining experiments used for training. This process is then repeated multiple times, specifically 10 in this case, to ensure a representative sampling of the OMM data and the development of statistics on the DMDc process.